A thin film with n = 1.6 coats a piece of glass. Destructive interference is observed for the reflection of light with wavelengths of 450 nm and 675 nm. What is the thinnest film for which this can occur? Express your answer in nm.
Answer #1


the condition for destructive interference is

2*n*t = m1*lamda1 = m2*lamda2

m1/m2 = lamda2/lamda1 = 675/450 = 1.5

let us take

m1 = 3 and m2 = 2

then

2*n*t = 3*lamda1 = 2*450*10^-9 = 900*10^-9

2*1.6*t = 900*10^-9

t = 281*10^-9 m = 281 nm